e) Assume that, due to the economic downturn, the president predicts that the corporation will earn only 4.5% per year on its money, not the previously anticipated 10% per year. What is the required amount of the annual revenue series over the 5-year period to be economically equivalent to the amount calculated in (d)?
1
Expert's answer
2020-05-12T10:53:55-0400

Let the corporation receive 5,000 each year from calculated in (d)

this is an annuity, find the future value of the annuity

"FV=A\times\frac{(1+r)^n-1}{r}"

"FV=5000\times\frac{(1+0.045)^5-1}{0.045}=27353.55"